OpenStudy (anonymous):

Given the function g(x) = -3x + 4, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these values. The value of g(-2) is smaller than the value of g(4) OR The value of g(-2) is the same as the value of g(4) OR The values of g(-2) and g(4) cannot be compared OR The value of g(-2) is larger than the vale of g(4)

OpenStudy (anonymous):

First, we solve for g(-2), meaning we plug in -2 for x. That gets us -3(-2)+4=6+4=10. Then, we solve for g(4). That gets us -3(4)+4=-12+4=-8. Then we compare.

OpenStudy (anonymous):

Awesome. Thank you!